Using the current income tax ordinance calculate the amount tax paid by an individual who is earning 50000 per month.


1
Expert's answer
2022-02-02T12:34:14-0500

Tax Payable

"0-9950\ @\ 10\%=(9950\times0.1)=950"

"9951-40,525\ @\ 12\%=" "(40,525-9950)\times0.12=3669"

"40,526-86,375\ @\ 22\%" "=(50,000-40,525)\times 0.22=2084.5"

"950+3,669+2,084.5=6,703.5"

Total tax payable = 6,703.5


Reference.

Foster, S. (2021, August 18). 2020-2021 tax brackets and federal income tax rates. Bankrate. https://www.bankrate.com/taxes/tax-brackets/